OpenStudy (anonymous):

The expressions x-6, 2x, and x^2 from the first three terms of a geometric progression. By calculating two different expressions for the common ratio, form and solve an equation in x to find possible values of the first term.

OpenStudy (anonymous):

for a,b,c to be in GP, b^2 = a*c we can use this to get 'x' from which the first term can be found. hope tht helps

OpenStudy (anonymous):

no sorry could you do some of the steps please

OpenStudy (anonymous):

did you know the above rule for terms in GP???

OpenStudy (anonymous):

what gp? and no

OpenStudy (anonymous):

geometric progression :P

OpenStudy (anonymous):

oh I see haha

OpenStudy (anonymous):

for a,b,c to be consecutive terms of a geometric progression, (a/b)=(b/c) -->(1) did you know this ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the same is rewritten by me as b^2 =a*c

OpenStudy (anonymous):

how do you simplify this though

OpenStudy (anonymous):